OpenStudy (anonymous):

solve by the substitution method 3m+n=23 m-6n=14

OpenStudy (anonymous):

so we solve the equation for one to substitute so we say n=23-3m then plug it into the other equation to solve for m m-6(23-3m)=14

OpenStudy (anonymous):

m-138+18m=14 19m=152 m=8

OpenStudy (anonymous):

then we can use any of the equations to find m so you can just substitute it in and let me know what you get

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and you can check it by substituting back the points